# Programming with Pixels: Computer-Use Meets Software Engineering

### v1

@@ 3.2 Key Features of Programming with Pixels
Future Adaptability.
As agents continue to evolve, PwP provides a unified environment for incorporating new benchmarks and training. For example, PwP is amenable to reinforcement learning due to its use of the standard gymnasium (Towers et al., 2024) interface and its checkpointing functionality. We show an example of how to interact with PwP in Figure 2. Checkpointing also allows for backtracking in search-based methods (Koh et al., 2024b; Putta et al., 2024). As software engineering practices progress and new IDE tools emerge, PwP incorporates them without additional engineering overhead. Further, as agents become more ubiquitous, it is imperative to evaluate their capabilities in pair-programming scenarios with human developers. PwP supports concurrent user-agent interaction, potentially enabling new kinds of pair programming or real-time collaboration studies. Finally, adding new tasks requires only minimal modification to configuration and evaluation files.
3.3 Infrastructure and Implementation
@@
As shown in 2, a Python API is provided for interaction, following a style similar to common reinforcement learning libraries such as gymnasium (Towers et al., 2024). The API abstracts away the complexity of container management, benchmark management, and handling observations and actions, allowing researchers to focus on agent development. Users can query the environment for the latest screenshot, issue an xdotool command, and receive updated states or rewards. The environment’s container configuration is flexible, allowing for software installations, customizable CPU/memory limits, and display settings (e.g., resolution). This versatility is crucial for large-scale evaluation, especially when tasks vary in complexity and resource needs.
4 PwP-Bench
We introduce PwP-Bench, a benchmark comprising 15 diverse software engineering tasks that span 8 programming languages and multiple modalities. Each task provides agents with access to the complete suite of tools available in the PwP environment. The purpose of PwP-Bench is to evaluate how well agents handle a broad range of software engineering (SWE) activities, thereby testing the generality of their code generation and SWE capabilities.
Tasks.
PwP-Bench contains 5400 instances covering 15 tasks, sourced from 13 existing code-generation datasets and 2 newly created by us. These tasks are designed to be representative of the breadth of software engineering—including tasks beyond conventional code generation to capture real-world complexities—and can be expanded as models excel in newer tasks. We followed three guiding principles: (1) tasks should require significant interaction with various SWE tools, (2) each task should necessitate multiple steps to complete, and (3) the overall benchmark should span multiple programming languages and modalities. Based on these principles, we collected diverse tasks and grouped them into four categories:
  • Code Generation and Editing: Evaluates the ability of agents to generate and edit code. This category includes datasets such as HumanEval for code completion, SWE-Bench (Jimenez et al., 2023) and SWE-Bench-Java (Zan et al., 2024) for resolving pull requests, DSBench (Jing et al., 2024) for data science tasks, and Res-Q (LaBash et al., 2024) or CanITEdit (Cassano et al., 2024) for code editing. Each dataset benefits from different tools. For example, SWE-Bench can take advantage of debuggers and linters, while DSBench may leverage an IPython kernel and tools for analyzing large data files. Code editing tasks can leverage refactoring utilities and repository searches, covering varied input-output formats and end goals.
  • Multimodal Code Synthesis: Involves creating code based on input images or other visual data. Examples include Design2Code (Si et al., 2024) for UI development, Chart2Mimic (Shi et al., 2024) for generating Python code from chart images, SWE-Bench-MM (Yang et al., 2024b) for multimodal code editing, and DSBench tasks that rely on images or PDF documents during data analysis.
  • Domain-Specific Programming: Focuses on specialized fields such as ethical hacking (CTF) (Yang et al., 2023a) and interactive theorem proving (miniCTX) (Hu et al., 2024). These tasks demand significant interactivity with IDE components. For example, theorem proving requires continuously inspecting states via the IDE, while CTF tasks often involve analyzing images, running executables, or installing VSCode extensions (e.g., hexcode readers).
  • IDE-Specific and General SWE Tasks: Recognizing that code generation is only one aspect of software engineering, we introduce two novel task sets to evaluate broader SWE skills. The first, IDE Configuration, evaluates an agent’s ability to modify IDE settings—such as themes, extension installations, and preferences—that are critical for effective tool use in a complex environment. The second, which we term General-SWE, targets non-code activities such as profiling, designing UI mockups, managing Kanban boards, and project refactoring. These tasks capture essential operational skills typically required by human developers but largely absent from conventional code generation benchmarks.
Figure 3 shows the distribution of tasks across all categories. An agent that succeeds across these tasks demonstrates strong potential for automating a wide range of software engineering activities. For instance, PwP-Bench covers Python, Java, JavaScript, HTML, CSS, Bash, SQL, and Lean, and requires agents to work with text, images, data files, and other data types. Furthermore, effective interaction with IDE tools is essential.
@@ 4 PwP-Bench
Figure 3: Distribution of tasks in PWP-Bench across four main categories: Code Generation and Editing, Multimodal Code Synthesis, Domain-Specific Programming, and General SWE Tasks. The inner ring shows the main categories while the outer ring shows specific datasets and tasks within each category.
@@
Benchmarking Design and Task Setup.
Every task is evaluated within the PwP environment. Unlike traditional benchmarks that provide structured, well-formatted context (for example, supplying all relevant schemas in text-to-SQL), PwP-Bench presents agents with an IDE containing a codebase rich in information. Specifically, an agent is provided with an initial environment state $S_{i}$ and an instruction $I$. The agent’s goal is to update the codebase to satisfy $I$ and transition to a final state $S_{f}$. Only $S_{f}$ is evaluated, using execution-based criteria (e.g., running unit tests). This setup requires agents to autonomously discover and extract relevant information from files, directories, and other resources—mirroring the challenges faced in real-world software development.
Many tasks in PwP-Bench require extensive multi-turn interactions and can be time-consuming. To support large-scale evaluations, PwP enables parallelized testing in a sandboxed environment, ensuring both security and reproducibility. Tasks can also be configured to restrict or partially allow internet access based on experimental needs.
Furthermore, as software engineering tasks evolve with advancements in model capabilities, our benchmark is designed to grow over time. New tasks can be incorporated by creating simple setup scripts that define the IDE’s initial state and evaluation logic, ensuring that PwP-Bench remains modular and adaptable.
PwP-Bench-Lite.
Because PwP-Bench contains more than 5400 instances in total, running a full evaluation can be computationally expensive. To address this, we also provide PwP-Bench-Lite—a smaller subset of 300 instances, made up of 20 random samples per task. This subset preserves the overall difficulty and distribution while ensuring equal representation for each task, thereby making rapid experimentation more accessible.
5 Agents in Programming with Pixels
The primary objective of Programming with Pixels is to enable general-purpose SWE agents. We evaluate state-of-the-art agents based on vision-language models in our environment. Each agent operates in a turn-based manner, receiving a screenshot each turn and returning an action (keyboard or mouse action) to progress toward the goal. This design is the same as used in previous works (Xie et al., 2024; Koh et al., 2024a) and we refer them to as computer-use agents. In practice, most vision-language models struggle with raw image inputs. To mitigate this, we incorporate Set-of-Marks (SoM) (Yang et al., 2023b), in which the agent receives both the raw image and a parse of available interface element (e.g., buttons, text fields). The agent then interacts with element IDs instead of raw pixel coordinates.
We evaluate two categories of computer-use agents. The first category only outputs keyboard and mouse clicks. That is, the agent uses the following observation and action space:
  • $\mathbf{O}$: a screenshot and set-of-marks annotations.
  • $\mathbf{A}$: keyboard and mouse clicks with set-of-marks.
The second category of computer-use agents has access to file and bash commands supplied by the environment through an API. Furthermore, a screenshot is received only when the screenshot action is called, instead of receiving it every turn. These actions are provided in PwP in a similar design principle as Anthropic computer-use (Anthropic, 2024) which consists of file operations such as ‘read file’, ‘create file’, and ‘string replace’. In summary, these agents use the following observation and action space:
  • O: a screenshot and set-of-marks annotations and text output from tools if used.
  • A: keyboard, mouse, and file and bash operation actions.
Finally, in Analysis 6.2, we create a tool-based agent design compatible with PwP. Specifically, we provide agents with domain-specific API calls that represent high-level actions (such as getting file structure) instead of UI interactions. Each high-level action is implemented as a sequence of low-level actions executed in PwP. This setting lets us test current state-of-the-art SWE agent patterns (e.g., Wang et al. (2024b)) within the PwP environment.
6 Experiments
Table 3: Performance Evaluation of Different Agents on PwP-Bench by Task Categories
Inputs Outputs Model Code Generation Multimodal Domain-Specific General Overall
& Editing Code Generation Code Generation SWE Tasks Avg
Screenshot + SoM Keyboard + Mouse GPT-4o 0.8% 12.4% 1.7% 10.0% 5.3%
GPT-4o-mini 0.8% 3.7% 0.0% 2.5% 1.7%
Gemini-Flash 0.0% 4.3% 0.0% 0.0% 1.2%
Gemini-Pro 2.5% 5.7% 0.0% 7.5% 3.%
Claude-Sonnet 10.0% 8.2% 5% 20.0% 10.2%
Screenshot + SoM + Tool Output Keyboard + Mouse + Tool Call (File, Bash) GPT-4o 37.0% 41.9% 28.3% 5.0% 32.3%
GPT-4o-mini 23.6% 17.5% 15.0% 5.0% 17.8%
Gemini-Flash 9.5% 11.7% 8.3% 2.5% 8.9%
Gemini-Pro 30.9% 16.7% 3.3% 5.0% 18.1%
Claude-Sonnet 52.1% 55.1% 43.3% 20.0% 46.8%
Experimental setup.
We evaluate two categories of baseline agents as described in Section 5. For each configuration, we test five state-of-the-art vision-language models (VLMs): Gemini-Flash-1.5, Gemini-Pro-1.5, GPT-4o, GPT-4o-mini, and Claude-3.5 Sonnet. With the exception of Claude—which is natively trained for UI interaction—the remaining models are provided with SoM.
At each timestep, an agent receives an observation (with the observation space determined by its category) and returns an action. The complete history of observations and actions is incorporated into the model’s context. For each task instance, the maximum number of iterations is capped at 20 steps; if the agent either exhausts these steps or issues a stop command, the environment’s final state is evaluated using task-specific metrics (see Appendix B for full details). Notably, the agent design remains the same throughout all tasks. Due to computational and budget constraints, we evaluate on PwP-Bench-Lite, which has 300 task instances.
6.1 Results
Figure 7: Example of Agent Hallucinating Screen Contents While the agent correctly mentions, search results are displayed (green text), it hallucinates an input field containing “disable import error” (red)
Table 3 summarizes performance across different agent architectures and base models over the four categories of PwP-Bench. As seen in the top half of the table, when using only primitive keyboard and mouse actions most agents have poor performance, with a maximum overall average of 10.2%. We attribute this poor performance primarily to limited visual grounding and an inability to interact effectively with the IDE—particularly for file editing and tool usage (see Section 6.2 for further analysis).
In contrast, when agents are granted access to file editing and bash operations through API calls rather than relying solely on UI interactions, we observe consistent improvements across all categories, with maximum average accuracy reaching 46.8%. Among the evaluated models, the Claude computer-use agent performs best, likely because it is specifically trained for UI interactions. We found that this agent is able to leverage basic IDE tools—such as HTML live preview, chart visualization, and file navigation—to boost performance on tasks requiring visual understanding and IDE navigation. Importantly, we find for the first time that a single computer-use agent can achieve performance comparable to and often surpassing (See Appendix C) state-of-the-art methods across a wide variety of software engineering (SWE) tasks—encompassing multiple languages, modalities, and domains—while operating within a single unified environment and interface, and no specific hand-crafted tools, unlike existing SoTA methods.
Nonetheless, as detailed in Section 6.2, the models struggle to fully take advantage of the IDE tooling. This is evidenced by the poor performance on the ‘General SWE’ dataset, where tasks are often as simple as editing IDE settings and often require fewer than four clicks to complete. As we show in Section 6.2, these tasks become simpler if the models could use the IDE tooling more effectively.
Overall, while the results point toward a promising direction for developing general computer-use SWE agents, significant improvements are still needed in visual grounding, tool usage, and planning. We analyze these next.
6.2 Analysis
Agents Demonstrate Poor Visual Grounding Capabilities.
Our qualitative analysis across multiple VLMs on PwP-Bench reveals significant limitations in visual understanding—even for basic IDE interactions. We identify two primary failure modes. First, models frequently fail to correctly identify the UI elements intended for interaction, as demonstrated in Figure 78. In agents using set-of-marks (SoM), this issue manifests as incorrect element selection, while without SoM it leads to inaccurate mouse positioning. Second, models struggle to comprehend the current UI state. As shown in Figure 910, they consistently fail to recognize highlighted elements, cannot detect linter errors indicated by wavy underlines (Figure 11), and often confuse active panels—resulting, for example, in typing into search bars rather than file editors (Figure 9). While similar issues have been documented in web and OS domains (Koh et al., 2024a; Xie et al., 2024), these limitations were primarily observed in models without UI-specific training. However, our work, shows even models explicitly trained for UI interaction (Anthropic, 2024), including Claude-Computer Use, exhibit these issues in PwP—likely due to the increased complexity of the IDE interface.
Agents Fail to Edit Files. File editing is a basic capability required in most SWE tasks. However, we find that the deficiencies in visual grounding significantly impact the file editing capabilities of current agents that use basic actions (clicking and typing). For example, even when provided with cursor location information in textual form, these models struggle to interpret such data amid complex UI elements. Models fine-tuned for UI interactions still commit basic editing errors—such as incorrect indentation and text misplacement—and are unable to recover from these errors (see Appendix for examples). We speculate these limitations could stem from two factors: (i) model overfitting to user interfaces in their training domains, or (ii) the increased complexity of the PwP IDE interface, which contains substantially more interactable elements than typical web or OS environments. Addressing these limitations represents an important direction for future work. Although direct file access via tool operations is available, UI-based editing confers unique advantages for tasks such as editing Jupyter notebooks, comparing changes, or modifying specific sections of large files. These results underscore two limitations: (i) current VLMs are challenged by complex UI interactions beyond simple web/OS interfaces (Xie et al., 2024; Koh et al., 2024a), and (ii) the inability to effectively perform UI-based editing prevents agents from leveraging valuable IDE features that could have improved their performance.
Figure 9: Example of Agent Misidentifying UI Elements The agent fails to identify the correct input field, typing ‘50’ into the settings search bar instead of the word wrap column setting field (red arrow).
Claude Computer-Use Agent Demonstrates Basic IDE Tool Proficiency.
Our analysis shows that the Claude Computer-Use agent successfully uses fundamental IDE functionalities, including file explorer navigation, file editing, search, browser-based live preview, and image generation and visualization capabilities. Figure 4 demonstrates the agent’s effective use of browser tools in UI replication tasks. Similarly, Figure 6 illustrates the agent’s ability to coordinate multiple tools while editing specific lines in a repository, relying solely on screenshot observations and primitive keyboard/mouse actions.
Second, we hypothesize that the agent has additional abilities to use tools that can be uncovered through prompting or fine-tuning. To investigate this, we considered the symbol renaming task in our ‘General-SWE’ benchmark, in which Claude initially achieves 0% accuracy when attempting the task. However, when explicitly instructed to use the renaming tool, its accuracy improves to 50% (see Appendix C).
Agents Struggle to Use IDE Functionality.
Despite these successes with basic IDE functionality, computer-use agents demonstrate significant limitations when interfacing with more complicated IDE tools. We observe no successful instances of debugger usage or symbol listing operations. Models without specific UI interaction training struggle even with fundamental tools such as HTML live preview, image visualization, and graph generation capabilities. While Claude demonstrates competency with basic tools, it fails to effectively use advanced tools like profilers or debuggers.
To further evaluate these capabilities, we developed the ‘General-SWE’ dataset, which focuses on software engineering activities (e.g., profiling, refactoring, debugging) that can be completed without direct code modification. Although these tasks sometimes require only 4-5 steps when using appropriate IDE tools, agents achieve minimal performance, highlighting substantial room for improvement.
Training Models to Use IDE Tools Better Would Improve Performance.
While a single computer-use agent design can perform well across a wide variety of tasks, our results indicate that these models do not fully exploit domain-specific tools. As an indication of the potential for performance gains if the agent was able to effectively use the IDE, we perform an “assisted” experiment.
For the assisted experiment, we manually engineer a set of API calls that are useful for the tasks. For example, in Design2Code, the assisted agents have an API call for a live HTML preview, while for SweBench it has API calls for retrieving repository structure and symbol outlines. Importantly, each API call is achievable using basic operations in the IDE, meaning that in principle an agent could learn to perform it. To ensure that each API call is achievable in the IDE, we implement each API call by executing a fixed sequence of low-level IDE actions, with the details abstracted away from the agent. A complete list of tools available in different environments is in Table 7 (see Appendix C).
Table 4 compares the performance of these assisted agents with that of standard computer-use agents across four datasets for which we manually created tools. The assisted agents achieved up to a 13.3% improvement in average scores relative to the non-assisted agents. This suggests that training agents to explore and use the built-in IDE functionality would yield performance gains. It also suggests that in the near term, we can get performance gains by introducing hand-engineered tools into the computer-use agent and incorporating existing agent designs in our PwP environment.
Table 4: Comparison of Different Agent Types Across Selected Tasks
SWE-Bench Design2Code Chartmimic BIRD (T2 SQL)
Primitive Agents 0% 23.5% 2.7% 0%
Computer Use 15% 48.1% 25.3% 7%
Assisted 19% 79.5% 61.6% 17%
Agents Are Incapable of Recovering from Errors. Next, we find that current agents show limited error recovery capabilities. When an action fails to execute correctly, models tend to persistently repeat the same failed action without exploring alternatives. Similarly, if an agent selects an incorrect action, it continues along an erroneous solution path without recognizing or correcting the mistake. In an experiment designed to probe this behavior, we deliberately suppressed one of the model’s actions. Despite the environment’s screenshot clearly showing an unchanged state, the models proceeded with their planned action sequence as though the suppressed action had succeeded. This behavior suggests a heavy reliance on memorized action sequences rather than dynamic responses to visual feedback, resulting in exponentially increasing errors and poor performance.
7 Conclusion
In this work, we introduce PwP, a unified environment that challenges the prevailing tool-based paradigm by enabling direct interaction with IDEs through basic computer-use actions like typing and clicking. This approach allows for a wide range of tasks to be modeled without language- or domain-specific modifications. Second, we introduce PwP-Bench, a unification of existing SWE datasets evaluated in PwP. We find that general-purpose computer-use agents can approach or outperform previous state-of-the-art results, without any task-specific modifications. This suggests that the dominant paradigm of building specialized text-based tools for SWE agents may be superseded by end-to-end computer-use agents. However, our analysis reveals that state-of-the-art agents are still incapable of using the extensive set of tools available in PwP, and could perform better if they could use them. Our work opens up an exciting new direction of development of computer-use agents for SWE tasks, an important step towards reaching truly general-purpose SWE agents. Finally, PwP provides a common platform that supports existing agents and benchmarks, making it a foundation for future research on SWE agents.
Acknowledgments.
We thank Daniel Fried, Graham Neubig, and Zora Wang, Saujas Vaduguru, Atharva Naik, Riyaz Ahuja, Weihua Du for helpful feedback, Google and Open AI credit programs, and Convergent Research.
Appendix A Programming with Pixels (PwP) Environment
A.1 Tools
Previous methods have proposed use of various hand-engineered tools. However, as shown in Table 5, PwP natively supports all these tools.
Method Hand-engineered Tools Supported in PwP
CodeAct (Wang et al., 2024a) File Edit, IPython, Bash
SWE-agent (Yang et al., 2024a) Search File, Search Text, File Edit
A.2 Comparison with Other Environments
In Table 1, we compare PwP with existing environments across multiple dimensions. We focus on environments designed for software engineering agents as well as those developed for web agents and operating system interaction. Since many prior code agent efforts do not explicitly separate their agent and environment architectures, we evaluate their environments based on their published interfaces (e.g., SWE-Bench, SWE-Bench-MM). We evaluate environments along the following dimensions:
  • Multi-modal support: The ability to process diverse input modalities including text, images, video, and audio
  • General action space: Whether the environment provides a general-purpose interface for tool interaction. For instance, OSWorld, WindowsAgentArena, and BrowserGym enable keyboard/mouse interactions through visual inputs, creating a general action and observation space. In contrast, environments like SWE-Bench and OpenHands restrict interactions to predefined tool-specific APIs. While OpenHands does allow BrowserGym interface, which in turns allows for general action and observation space, the generality is limited to browser interactions and does not apply to other SWE-related tools.
  • Observation space: The range of environmental states and feedback accessible to an operating agent
  • State Checkpointing: Native support for preserving file system and process states, a capability unique to PwP
  • Tools: The set of available tools and capabilities accessible to agents
While some environments offer general action and observation spaces, they are not specifically designed for software engineering tasks. Programming with Pixels uniquely combines SWE-specific design with general-purpose interaction capabilities, enabling the development of general-purpose computer-use agents for software engineering.
Appendix B PwP-Bench
Metrics
Appendix C Results
Table 6 presents comprehensive results for all agent designs across 15 datasets in PwP-Bench. For comparison, we include previously reported state-of-the-art results, which represent the best performance achieved on individual datasets through various specialized approaches including finetuned models, custom tool interfaces, specific pipelines, prompts, inference strategies, and verifiers.
It is important to note that direct comparisons on individual datasets may not provide a complete picture, as state-of-the-art results often utilize different model versions, computational budgets, and domain-specific tools. For example, the 55% performance on SWE-Bench was achieved using significantly longer agent trajectories compared to PwP’s 20-iteration limit. Given statistical variations and differing evaluation conditions, we emphasize evaluating agents across the full range of tasks rather than focusing on individual dataset performance, as this better aligns with our goal of developing general-purpose SWE agents. We make our best effort to include the latest publicly available results, however, there may be minor discrepancies.
Notably, on the Intercode dataset (Capture the Flag), the computer-use agent significantly outperforms the state-of-the-art method despite the latter (SWE-Agent-Enigma) employing numerous specialized tools for static analysis, dynamic analysis, and networking (See Table 5). This improvement is statistically significant (p-value = 0.014, McNemar’s test). While the agent did not rely on specific IDE tooling, this result highlights the advantages of a simple, general agent design and suggests the potential benefits of allowing agents to automatically identify optimal tools rather than imposing human priors on the system.
Table 6: Performance Evaluation of Different Models Across Task Categories. Leged: HE: HumanEval, SB: SWEBench, SJ: Swebench-Java, RQ: ResQ, CI: CaniteEdit, ST: SWTBench, DC: Design2Code, CM: ChartMimic, DS: DSBench, SM: Swebench-MM, IC: Intercode-CTF, BD: Bird SQL, MC: Minictx, VS: VSCode, GS: No-Code SWE Tasks. *Much more costly method, ** Estimated Value since only Pass@8 is reported.

### v2

@@ A.2 Comparison with Other Environments
A.3 Infrastructure and Implementation
@@
As shown in Figure 9, a Python API is provided for interaction, following a style similar to common reinforcement learning libraries such as gymnasium (Towers et al., 2024). The API abstracts away the complexity of container management, benchmark management, and handling observations and actions, allowing researchers to focus on agent development. Users can query the environment for the latest screenshot, issue an xdotool command, and receive updated states or rewards. Examples of xdotool commands include ‘xdotool mousemove 1000 1200’ and ‘xdotool type ’hello world” and are shown in Figure 9. The environment’s container configuration is flexible, allowing for software installations, customizable CPU/memory limits, and display settings (e.g., resolution). This versatility is crucial for large-scale evaluation, especially when tasks vary in complexity and resource needs. Finally, the environment has been tested on three different operating systems: Ubuntu, MacOS, and Windows.
Appendix B PwP-Bench
@@
Figure 10: Distribution of tasks in PWP-Bench across four main categories: Code Generation and Editing, Multimodal Code Synthesis, Domain-Specific Programming, and General SWE Tasks. The inner ring shows the main categories while the outer ring shows specific datasets and tasks within each category. Note that the figure is not drawn based on relative size of tasks.
Table 4: Number of instances for each task in PwP-Bench
Task Number of Instances Languages
HumanEval 165 Python
Design2Code 485 HTML/CSS/JS
ChartMimic 600 Python
InterCode 100 Python, Bash
RES-Q 100 Python
CanItEdit 105 Python
VSCode 20 -
Bird 500 SQL
DSBench 112 Python
SWE-bench 2000 Python
SWE-Bench-Multilingual 91 C++, Typescript, Javascript, Rust, Go, C, Ruby, PhP, Java
Swebench-$M$ 510 Javascript
SWT-Bench 276 Python
Minictx 381 Lean
General SWE 20 -
B.1 Tasks
Figure 10 shows the set of tasks across all categories. Further, Table 4 shows the number of instances for each task in the full benchmark, along with the languages used in each of the tasks. PwP-Bench-Lite contains 300 instances, which is a random sample of 20 instances from each task.
B.2 Evaluation
All tasks are evaluated using programmatic verifiers. These verifiers are typically run on an separate environment, not accessible to the agent. This typically works based on fetching relevant files and information from the agent environment, and then running through task-specific evaluation scripts on a separate environment. However, to the user, this is abstracted away, and they simply have to call ‘env.get_reward()’ to get the exact score or correctness signal based on task.
Metrics
VSCode and General SWE Tasks
In this section, we detail the VSCode and General SWE tasks in PwP-Bench, created by us. The VSCode tasks are mostly designed to evaluate the ability of agents to use basic VSCode features, such as renaming all instance of a symbol in file, installing extensions, changing themes, modifying specific settings. All these tasks are evaluated based on final IDE state, either by invoking the ‘code’ cli tool, configuration files stored in environment filesystem, or through direct access to VSCode state provided by PwP (see  subsection A.3). General-SWE tasks, involves 5 categories of tasks: 1.) QA based on code profiling (evaluated based on final answer by model which requires using appropriate profiling tools), 2.) code refactoring (assessed through automated tests on the final repository state), 3.) debugging bugs in standard libraries (evaluated based on the correctness of final code state), 4.) UI mockup design (assessed using CLIP scores), and 5.) code restoration, where the agent leverages VSCode’s timeline feature to recover corrupted codebases, evaluated by the correctness of the restored state.
Comparison with Other Benchmarks
In  Table 5, we further compare PwP-Bench with other existing benchmarks.
@@ B.2 Evaluation
Appendix C Related Work
C.1 Comparison to Software Engineering Agents
Task-specific SWE benchmarks
Early neural code generation approaches were typically evaluated on fixed input-output pairs—for example, generating code from docstrings (Chen et al., 2021) or from general textual descriptions (Austin et al., 2021). Subsequent benchmarks extended these evaluations to interactive settings, such as resolving GitHub pull requests or writing unit tests for real-world code repositories (Jimenez et al., 2023; Zan et al., 2024; Mündler et al., 2025). More recently, efforts have broadened the scope of code generation to include multimodal tasks, where vision models must interpret images to generate correct code or edits (Si et al., 2024b; Shi et al., 2024; Jing et al., 2024; Yang et al., 2024b). However, each of these benchmarks is confined to specific languages, modalities, or task types. In contrast, our proposed PwP-Bench unifies these diverse evaluations into a single framework, encompassing multimodal and multilingual challenges that require interaction with a broad suite of IDE tools. Using this unified approach we reproduce the performance of established benchmarks and encourage the development of general-purpose agents capable of handling a variety of new software engineering tasks. We further compare our work with previous efforts in Tables 3 and 5.
Software Engineering (SWE) Agents
C.2 Comparison to General Visual and Computer-Use Agents
Visual Agents and Computer-Use Agents
A family of recent multimodal agent benchmarks require agents to operate user interfaces using a predefined, limited set of actions (e.g., new_tab, go_back, click [element id](Koh et al., 2024; Deng et al., 2023; Zheng et al., 2024) . These visual agents typically rely on additional prompting—such as set-of-marks techniques that supply an HTML accessibility tree containing textual and positional information—to overcome their inherent poor visual grounding capabilities (Yang et al., 2023a). Despite such aids, these agents often fail when faced with the complex and dense IDE interfaces found in our environment.
Expressive Agent Environments
Prior work on expressive agent environments has predominantly targeted the web domain (Koh et al., 2024; Deng et al., 2023), entire operating systems (Xie et al., 2024; Bonatti et al., 2024; Rawles et al., 2023), or other general scenarios (Xu et al., 2024). Some of these environments, such as OSWorld (Xie et al., 2024), feature general action and observation spaces similar to ours. However, although these benchmarks are capable of expressing a wide range of tasks, they do not focus on the unique challenges inherent to software engineering within an IDE. For example, while OSWorld offers a broad set of tasks, it is not specifically designed for SWE, resulting in increased computational overhead. Software engineering is a diverse and important domain that merits its own dedicated environment.
Appendix D Experimental Setup and Implementation Details
D.1 Agent Design
In addition to the details mentioned in Section 5, we provide more implementation details in this section. First, the exact version numbers used for different API models are: gpt-4o-2024-11-20, gpt-4o-mini-2024-07-18, claude-3-5-sonnet-20241022, gemini-1.5-flash-preview-001, gemini-1.5-pro-preview-001, claude-3-7-sonnet-20250219, claude-sonnet-4-20250514. For the three Claude models, we use the computer-use variants by passing the ‘computer-use’ beta flag in API calls. For open-weights models, we run inference on 8 L40s using vLLM. We use temperature=0.3 consistently across models. For our main experiments, the number of iterations is set to 20 because: a.) for most tasks, 20 iterations is enough to complete the task, b.) increasing the number to more than 20 would increase the computational cost, and since some models didn’t support caching at the time of running the experiments, the cost grows quadratically, c.) we ran experiments with 250 steps on Claude-Sonnet-4.0 on SWE-bench related datasets (see Appendix F); however, we found no difference in trends.
D.2 Mini-SweAgent
For the mini-sweagent, we use Claude-4.0 Sonnet. We use the same code as the official source code (SWE-agent, 2024), except that we modify it for multimodal tasks so that the agent receives required images as input in its prompt.
Appendix E Results
Table 6 presents comprehensive results for all agent designs across 15 datasets in PwP-Bench.
E.1 Comparison with best reported Specialized SWE Agents
In this section, we compare computer-use agents with the best reported specialized SWE agents scores on individual datasets. In particular, for each dataset, we use 3 different strategies to identify the best reported scores:
  • Citations: For each dataset, we manually go through the citations and find the most relevant works and look for reported scores.
  • Official Leaderboard: For some datasets, such as SWE-Bench, we use the official leaderboard to find the best reported scores.
  • Web-Search Agents: We further prompt ChatGPT-5 thinking to find the latest and highest reported scores on each of the datasets. We then manually verify the results based on the links provided.
For each dataset, we follow all three strategies and take the highest reported score. Typically these results are achieved using specialized approaches including finetuned models, custom tool interfaces, specific pipelines, prompts, inference scaling, and verifiers. Therefore, it is important to note that direct comparisons on individual datasets may not provide a complete picture. Further, since our evaluations are done on 20 examples from the whole dataset, the results may not be directly comparable. Further, while we make our best effort to include the latest publicly available results, there may be still be discrepancies. Finally, the search was conducted on 22nd September 2025, and future numbers may change.
We now list the best reported scores for each dataset:
  • HumanEval: QualityFlow (Hu et al., 2025) achieves 98.8% performance using Claude-3.5-Sonnet.
  • SWE-Bench: Highest scores (75.2%) are achieved by a method named TRAE agent (Team et al., 2025), with best reported performance with Claude-4-Sonnet as base model as 74.6%.
  • SWE-Bench-Multilingual: Highest score publicly reported is 43% (Yang et al., 2025) using Claude-3.7-Sonnet and Swe-agent framework (Yang et al., 2024a).
  • ResQ: Highest score publicly reported is 58% (LaBash et al., 2024) using Claude-3.5-Sonnet in the official dataset report.
  • SWT-Bench: Highest score publicly reported is 63.3% (Cassano et al., 2024) using GPT-4o in the official dataset report.
  • Design2Code: Highest score publicly reported is 90.2% (Si et al., 2024a) using Claude-3.5-Sonnet in the official dataset report.
  • Chartmimic: Highest score publicly reported is 86.46% using GPT-4o and METAL method (Li et al., 2025). Further they use inference scaling with n=5.
  • Intercode-CTF: The publicly reported state of the art number is 72% using SWE-Agent-Enigma (Abramovich et al., 2024). This is much smaller than the numbers reported by our computer-use agent evaluation, which reaches 100% with the same Claude-3.5-Sonnet model. This is surprising, since the method employed numerous specialized tools for static analysis, dynamic analysis, and networking, and we confirmed that the improvement is statistically significant (p-value = 0.014, McNemar’s test).
  • BIRD: The best reported score is 76.14% (Shkapenyuk et al., 2025) as per the numbers reported in offciail leaderboard.
  • SWE-Bench-Multimodal: The best reported score is 35.98% using scaffolding over O3, and 34.33% when using OpenHands-Versa (Soni et al., 2025) with Claude-4-Sonnet.
Overall, the results are often much higher than the numbers achieved by computer-use agents, even with access file and bash APIs. Overall, the discussion points out that at present specialized software-engineering agents still perform better, and built scaffolding around computer-use agents might also be helpful.
Table 6: Performance Evaluation of Different Models Across Task Categories. Leged: HE: HumanEval, SB: SWEBench, SJ: Swebench-Multilingual, RQ: ResQ, CI: CaniteEdit, ST: SWTBench, DC: Design2Code, CM: ChartMimic, DS: DSBench, SM: Swebench-$M$, IC: Intercode-CTF, BD: Bird SQL, MC: Minictx, VS: VSCode, GS: General-SWE Tasks.